# De-Confusing Pseudo-Labels in Source-Free Domain Adaptation

### v2

@@ Abstract
Source-free domain adaptation (SFDA) aims to adapt a source-trained model to an unlabeled target domain without access to the source data. SFDA has attracted growing attention in recent years, where existing approaches focus on self-training that usually includes pseudo-labeling techniques. In this paper, we introduce a novel noise-learning approach tailored to address noise distribution in domain adaptation settings and learn to de-confuse the pseudo-labels. More specifically, we learn a noise transition matrix of the pseudo-labels to capture the label corruption of each class and learn the underlying true label distribution. Estimating the noise transition matrix enables a better true class-posterior estimation, resulting in better prediction accuracy. We demonstrate the effectiveness of our approach when combined with several SFDA methods: SHOT, SHOT++, and AaD. We obtain state-of-the-art results on three domain adaptation datasets: VisDA, DomainNet, and OfficeHome.
@@ 1 Introduction
In SFDA, a key obstacle is to reduce the accumulation of errors caused by domain misalignment. Existing methods in SFDA focus on self-training using target pseudo-labels and entropy minimization techniques. In [27] the effective reduction of error accumulation was achieved by assigning pseudo-labels to the target data using class clusters in the penultimate layer of the model, where the pseudo-labels were determined based on distances from the centroids of the clusters. However, due to the domain shift, these generated pseudo-labels tend to be noisy. For this reason, several methods have focus on the refinement of the target pseudo-labels during training (e.g., [4, 17, 59, 61]). One common solution is to update the pseudo-labels at each epoch, thus refining them to better match the target distribution [27, 61]. Recently, in [59] it was argued that this approach leads to memorizing the label noise. Therefore, they suggested using an early-learning regularization so the pseudo-labels will not undergo drastic changes during the target adaptation process.
In this work we do not refine the pseudo-labels during target training. Instead, we approach SFDA by adopting a learning with label-noise (LLN) perspective and suggest a novel approach tailored for domain adaptation settings to De-Confuse the Pseudo-Labels. We call our method DCPL. It involves modifying the established technique of optimizing a noise transition matrix (or a confusion matrix) to adapt the classifier predictions to the pseudo-labels [44, 47]. Traditional learning-with-label-noise (LLN) methods typically address label noise stemming from errors made by human annotators during labeling. In our specific scenario, the label noise is derived from pseudo-labels generated with a specified source model. This label noise emerges from domain shifts, where the characteristics of data in the target domain deviate from those in the source domain used for training. In this context, the outputs of the source model carry valuable information about the noise distribution. This information can be leveraged to improve the estimation of the transition matrix. We propose integrating this knowledge into the noise transition matrix learning method to achieve a more accurate prediction of the underlying true label distribution, consequently enhancing the accuracy of classifier predictions. Empirically, we demonstrate that our approach is easy to integrate with current source-free unsupervised domain adaptation (SFDA) techniques and yields improved prediction accuracy compared to baseline methods, hence achieving new state-of-the-art results. Our main contributions are as follows:
@@
We demonstrate the effectiveness of our approach by implementing it across several SFDA methods: SHOT, SHOT++, and AaD. Our results establish new state-of-the-art performance on three domain adaptation datasets: VisDA, DomainNet, and OfficeHome.
@@ 2 Related Work
Unsupervised domain adaptation (UDA) attracted much attention in recent years [50, 8, 9, 49, 42]. One common approach for addressing UDA involves employing distributional matching (DM). This is usually done by directly minimizing the domain discrepancy statistics [45, 34] or by using some form of domain-adversarial learning [9, 49]. However, in [23] it was shown that DM-based approaches have limited generalization ability under what they term “realistic shifts". One such shift is when the source label distribution is balanced while the target label distribution is long-tailed. A possible remedy proposed in [23] is to use self-supervised methods which were found to be useful for UDA [46, 10, 55, 1].
@@
Recently, in light of the growing emphasis on privacy and data protection, there has been a rising interest in source-free domain adaptation [24, 27, 56, 52, 40, 39, 15]. In SFDA, a model is first trained on the source data and then updated based on the target data, usually using unsupervised techniques and self-training due to the absence of labeled data [27]. Some studies modify the training procedure of the source model to ease the adaptation process by using self-supervision techniques [20], mixup augmentation [21], and/or uncertainty quantification [41]. Another line of research adds unsupervised loss terms to encourage discriminability and diversity in the output during the target adaptation process [27, 5, 57, 58, 56]. For instance, AaD [56] designs the loss such that it encourages the prediction of neighboring samples to be similar and the prediction of non-neighboring samples to be different. This results in simple discriminability and diversity terms. In SHOT [27] the authors use the conditional entropy and marginal entropy of the predictions as well as cross-entropy with target pseudo-labels. Using target pseudo-labels, the authors managed to reduce the error accumulation caused by domain misalignment. SHOT++ [28] builds on SHOT by adding a rotation-prediction head to allow self-supervised representation learning on the target domain. The samples are then divided into unlabeled and "labeled" based on the prediction confidence, and an additional step of semi-supervised learning is applied [3].
@@
Several SFDA studies have focused on correcting the pseudo-labels of the target samples under the assumption that they are noisy [4, 17, 19, 30, 59]. In [17] only reliable samples from the target domain were chosen based on prediction statistics. However, this may lead to under-representation of the target domain data distribution. In [19] the authors used label refinement based on similarity statistics between classes across consecutive epochs. However, in cases of high fluctuations in the class assignment, the similarity matrix may be non-diagonally dominant resulting in uninformative or even damaging labels. In [30] the labels were refined based on the predictions of neighboring examples. However, it requires storing a memory bank of examples which may impose a burden in low-resource environments. Yi et al. [59] noticed that the early-time training phenomenon is present in the SFDA settings. Hence, they added a regularization term to maintain the generated pseudo-labels close to their initial values, thus preventing significant changes during training. Recently, a co-learning strategy [61] was suggested to improve the quality of the generated target pseudo-labels using robust pre-trained networks such as Swin-B [33]. In addition, they refined the pseudo-labels and replaced uncertain ones with the current model output.
@@
In general, methods for LLN assume either instance-dependent noise or instance-independent noise [43]. In this study, we assume that the true label is conditionally independent of the features given the observed label. Common approaches for handling label noise with deep learning models include adding regularization [38, 31, 53], modifying the loss functions and values [32, 11, 35, 2], sample selection [13, 16, 60, 26, 48], and architectural changes, such as adding a noise adaptation layer [44, 54, 12]. Our approach belongs to the last category since it is simple to incorporate into existing learning systems, easily scaled, and tends to work well. One obstacle to learning a confusion matrix is to prevent converging to degenerate solutions. Various methods have been proposed to address this issue. In [44, 47] it was suggested to regularize the matrix trace. Goldberger and Ben-Reuven [12] uses a joint learning of the network and the confusion matrix with a careful matrix initialization. In [62] the authors suggest maximizing the total variation between the prediction vectors of samples. Li et al. [25] proposed controlling the volume of the simplex formed by the columns of the confusion matrix by minimizing its log determinant. Lin et al. [29] advocated a bi-level optimization process of the network’s parameters and the confusion matrix. In this study, we chose to minimize the trace as we found it to work well.
@@ 4.2 Adaptation as learning with noisy pseudo-labels
where $\bar{p}_{ji}\triangleq\mathbb{E}_{p(x|y=i)}[p(\arg\max\limits_{k}{p_{\theta}}^{(k)}(y|x)=j|y=i)]$.
Moreover, it was shown in [47] (Lemma 1) that obtaining $\hat{T}$ with a minimal trace uniquely coincides with $T$. Note that assumption (2) holds by bringing $\hat{T}\cdot{p_{\theta}}(y|x_{t})$ closer to the true noisy distribution $\tilde{y}$, which is encouraged by minimizing the cross-entropy. As for assumption (1), using a large pre-trained network (Swin-B [33]) and a source model that is trained on a source domain with a reasonable domain gap from the target domain will probably generate a diagonally dominant $T$ (see Section 5 for a transition matrix example).
@@ 5.2 Implementation Details
For the target adaptation process, we generate the pseudo-labels at the beginning of training using a pre-trained network Swin-Base [33], since Swin transformers [33] were shown to be robust under domain shifts [18]. In our experiments, we used two pre-trained SwinB networks, one of which was pre-trained on Imagenet-1K, and the other which was pre-trained on Imagenet-22K and then finetuned on Imagenet-1K (ImageNet-22K-1K). The learning rate is set to 1e-2 for OfficeHome, and 1e-3 for VisDA and DomainNet. The temperature $\tau$ for calculating $T_{c}$ is set to 0.01 when using the Imagenet-1K pretrained network and 0.05 for ImageNet-22K-1K. $\gamma$ is set to 1 for VisDA and to 0.01 for DomainNet and OfficeHome, to counter the larger overall magnitude of the prior regularization term due to the increased number of classes. $\lambda$ was set to 0.01 for all datasets. We train the adapted networks for 50 epochs in all cases. All of our results report the mean accuracy over three runs with different random seeds (2019, 2020, and 2021) as was done in [27].
@@ 5.3 Main Results
Table 4: Prior loss contribution
@@
Table 5: Influence of PL network
@@ 5.4 Ablation Study
Table 6 presents a performance comparison of DCPL to an oracle implementation on the DomainNet dataset. The oracle version was implemented using the real noise transition matrix based on the target labels (which we do not have access to in UDA). The comparison to the oracle allows us to analyze the performance increase potential of DCPL when the estimated transition matrix is made accurate. In the table, we also compare DCPL to a naive baseline implementation in which the transition matrix is set to a fixed identity matrix ($\hat{T}=I$); i.e., if we assume noiseless pseudo-labels. Comparing this naive baseline to the oracle version of DCPL shows an increase of 7.2% to 72.5% and 5.3% to 76.8%, using SwinB-1K and SwinB-22K-1K respectively. As expected, DCPL achieves an intermediate accuracy of 69.6% and 73.7%. These results show that better estimations of the transition matrix can generate significant improvement gains.

### v3

@@ Abstract
Source-free domain adaptation aims to adapt a source-trained model to an unlabeled target domain without access to the source data. It has attracted growing attention in recent years, where existing approaches focus on self-training that usually includes pseudo-labeling techniques. In this paper, we introduce a novel noise-learning approach tailored to address noise distribution in domain adaptation settings and learn to de-confuse the pseudo-labels. More specifically, we learn a noise transition matrix of the pseudo-labels to capture the label corruption of each class and learn the underlying true label distribution. Estimating the noise transition matrix enables a better true class-posterior estimation, resulting in better prediction accuracy. We demonstrate the effectiveness of our approach when combined with several source-free domain adaptation methods: SHOT, SHOT++, and AaD. We obtain state-of-the-art results on three domain adaptation datasets: VisDA, DomainNet, and OfficeHome.
@@ 1 Introduction
In SFDA, a key obstacle is to reduce the accumulation of errors caused by domain misalignment. Existing methods in SFDA focus on self-training using target pseudo-labels and entropy minimization techniques. In [28] the effective reduction of error accumulation was achieved by assigning pseudo-labels to the target data using class clusters in the penultimate layer of the model, where the pseudo-labels were determined based on distances from the centroids of the clusters. However, due to the domain shift, these generated pseudo-labels tend to be noisy. For this reason, several methods have focused on the refinement of the target pseudo-labels during training (e.g., [5, 18, 61, 63]). One common solution is to update the pseudo-labels at each epoch, thus refining them to better match the target distribution [28, 63]. In [63], the authors handle this problem by pseudo-labeling using a strong pre-trained network and filtering out samples with low confidence.
In this work we do not refine the pseudo-labels during target training and do not make any explicit decision on the PLs correctness. Instead, we approach SFDA by adopting a learning with label-noise (LLN) perspective and suggest a novel approach tailored for domain adaptation settings to De-Confuse the Pseudo-Labels. We call our method DCPL. It involves modifying the established technique of optimizing a noise transition matrix (or a confusion matrix) to adapt the classifier predictions to the pseudo-labels [45, 49]. Traditional learning-with-label-noise (LLN) methods typically address label noise stemming from errors made by human annotators during labeling. In our specific scenario, the label noise is derived from pseudo-labels generated with a specified source model. This label noise emerges from domain shifts, where the characteristics of data in the target domain deviate from those in the source domain used for training. In this context, the outputs of the source model carry valuable information about the noise distribution. This information can be leveraged to improve the estimation of the transition matrix. We propose integrating this knowledge into the noise transition matrix learning method to achieve a more accurate prediction of the underlying true label distribution, consequently enhancing the accuracy of classifier predictions. Empirically, we demonstrate that our approach is easy to integrate with current source-free unsupervised domain adaptation (SFDA) techniques and yields improved prediction accuracy compared to baseline methods, hence achieving new state-of-the-art results. Our main contributions are as follows:
@@
We demonstrate the effectiveness of our approach by implementing it across several SFDA methods: SHOT, SHOT++, and AaD. Our results establish new state-of-the-art performance on three domain adaptation datasets: VisDA, DomainNet, and OfficeHome. Our code can be found at: https://github.com/ssi-research/DCPL_SFDA.
@@ 2 Related Work
UDA attracted much attention in recent years [52, 9, 10, 51, 43]. One common approach for addressing UDA involves employing distributional matching (DM). This is usually done by directly minimizing the domain discrepancy statistics [46, 35] or by using some form of domain-adversarial learning [10, 51]. However, in [24] it was shown that DM-based approaches have limited generalization ability under what they term “realistic shifts". One such shift is when the source label distribution is balanced while the target label distribution is long-tailed. A possible remedy proposed in [24] is to use self-supervised methods which were found to be useful for UDA [47, 11, 57, 1].
@@
Recently, in light of the growing emphasis on privacy and data protection, there has been a rising interest in source-free domain adaptation [25, 28, 59, 54, 41, 40, 16]. In SFDA, a model is first trained on the source data and then updated based on the target data, usually using unsupervised techniques and self-training due to the absence of labeled data [28]. Some studies modify the training procedure of the source model to ease the adaptation process by using self-supervision techniques [21], mixup augmentation [22], and/or uncertainty quantification [42]. Another line of research adds unsupervised loss terms to encourage discriminability and diversity in the output during the target adaptation process [28, 6, 58, 60, 59]. For instance, AaD [59] designs the loss such that it encourages the prediction of neighboring samples to be similar and the prediction of non-neighboring samples to be different. This results in simple discriminability and diversity terms. In SHOT [28] the authors use the conditional entropy and marginal entropy of the predictions as well as cross-entropy with target pseudo-labels. Using target pseudo-labels, the authors managed to reduce the error accumulation caused by domain misalignment. SHOT++ [29] builds on SHOT by adding a rotation-prediction head to allow self-supervised representation learning on the target domain. The samples are then divided into unlabeled and "labeled" based on the prediction confidence, and an additional step of semi-supervised learning is applied [4]. The authors of [61] noticed that the early-time training phenomenon is present in the SFDA settings. Hence, they added an auxiliary loss to prevent the model from deviating too much from early-time predictions.
@@
Several SFDA studies have focused on correcting the pseudo-labels of the target samples under the assumption that they are noisy [5, 18, 20, 31, 61, 2]. In [18] only reliable samples from the target domain were chosen based on prediction statistics. However, this may lead to under-representation of the target domain data distribution. In [20] the authors used label refinement based on similarity statistics between classes across consecutive epochs. However, in cases of high fluctuations in the class assignment, the similarity matrix may be non-diagonally dominant resulting in uninformative or even damaging labels. In [31] the labels were refined based on the predictions of neighboring examples. However, it requires storing a memory bank of examples which may impose a burden in low-resource environments. Recently, a co-learning strategy [63] was suggested to improve the quality of the generated target pseudo-labels using robust pre-trained networks such as Swin-B [34]. In addition, they refined the pseudo-labels and replaced uncertain ones with the current model output.
@@
In general, methods for LLN assume either instance-dependent noise or instance-independent noise [44]. In this study, we assume that the true label is conditionally independent of the features given the observed label. Common approaches for handling label noise with deep learning models include adding regularization [39, 32, 55], modifying the loss functions and values [33, 12, 36, 3], sample selection [14, 17, 62, 27, 50], and architectural changes, such as adding a noise adaptation layer [45, 56, 13]. Our approach belongs to the last category since it is simple to incorporate into existing learning systems, easily scaled, and tends to work well. One obstacle to learning a confusion matrix is to prevent converging to degenerate solutions. Various methods have been proposed to address this issue. In [45, 49] it was suggested to regularize the matrix trace. The authors of [13] use a joint learning of the network and the confusion matrix with a careful matrix initialization. In [64] it was suggested to maximize the total variation between the prediction vectors of samples. In [26] the authors proposed controlling the volume of the simplex formed by the columns of the confusion matrix by minimizing its log determinant. The authors of [30] advocated a bi-level optimization process of the network’s parameters and the confusion matrix. In this study, we chose to minimize the trace as we found it to work well.
@@ 4.2 Adaptation as learning with noisy pseudo-labels
where $\bar{p}_{ji}\triangleq\mathbb{E}_{p(x|y=i)}[p(\arg\max\limits_{k}{p_{\theta}}^{(k)}(y|x)=j|y=i)]$. Moreover, it was shown in [49] (Lemma 1) that obtaining $\hat{T}$ with a minimal trace uniquely coincides with $T$. Note that assumption (2) holds by bringing $\hat{T}\cdot{p_{\theta}}(y|x_{t})$ closer to the true noisy distribution $\tilde{y}$, which is encouraged by minimizing the cross-entropy. As for assumption (1), using a large pre-trained network (Swin-B [34]) and a source model that is trained on a source domain with a reasonable domain gap from the target domain will probably generate a diagonally dominant $T$ (see Section 5 for a transition matrix example).
@@ 5.2 Implementation Details
For the target adaptation process, we generate the pseudo-labels at the beginning of training using a pre-trained network Swin-Base [34], since Swin transformers [34] were shown to be robust under domain shifts [19]. In our experiments, we used two pre-trained SwinB networks, one of which was pre-trained on Imagenet-1K, and the other which was pre-trained on Imagenet-22K and then finetuned on Imagenet-1K (ImageNet-22K-1K). The learning rate is set to 1e-2 for OfficeHome, and 1e-3 for VisDA and DomainNet. The temperature $\tau$ for calculating $T_{c}$ is set to 0.01 when using the Imagenet-1K pretrained network and 0.05 for ImageNet-22K-1K. $\lambda$ was set to 0.01 [49] for all datasets to avoid overfitting and maintain simplicity for the practitioners. For $\gamma$ we chose 0.01 for OfficeHome and DomainNet which have a large number of classes and 1 for VisDA with a smaller amount of classes to account for the ratio in class amount (squared due to square in the loss term). We train the adapted networks for 50 epochs in all cases. All of our results report the mean accuracy over three runs with different random seeds (2019, 2020, and 2021) as was done in [28].
@@ 5.3 Main Results
Table 4: Prior loss (Eq. 8) contribution to DCPL performance
@@
Table 5: Influence of PL network (SwinB-1K versus SwinB-22K-1K) on performance
@@ 5.4 Ablation Study
Table 6 presents a performance comparison of DCPL to an oracle implementation on the DomainNet dataset. The oracle version was implemented using the real noise transition matrix based on the target labels (which we do not have access to in UDA). The comparison to the oracle allows us to analyze the performance increase potential of DCPL when the estimated transition matrix is made accurate. In the table, we also compare DCPL to a naive baseline implementation in which the transition matrix is set to a fixed identity matrix ($\hat{T}=I$); i.e., if we assume noiseless pseudo-labels and do not learn the TM. Comparing this naive baseline to DCPL shows an increase of 4.3% using SwinB-1K and 2.2% using SwinB-22K-1K. This demonstrates the contribution of TM over SwinB. Comparing this naive baseline to the oracle version of DCPL shows an increase of 7.2% to 72.5% and 5.3% to 76.8%, using SwinB-1K and SwinB-22K-1K respectively. As expected, DCPL achieves an intermediate accuracy of 69.6% and 73.7%. These results show that better estimations of the transition matrix can generate significant improvement gains.